OpenStudy (anonymous):

Phil bought 12 toy cars. Of these cars, 2 are red, 4 are white, 3 are yellow, and 3 are green. If Phil puts them all in a bag, what is the probability he will blindly pull out a white toy car? 33.3% 25% 42% 16.7%

jimthompson5910 (jim_thompson5910):

P(White) = (# of white)/(# total) P(White) = 4/12 P(White) = 1/3 So the probability is 1/3 which as a percentage is 33.3%

OpenStudy (anonymous):

thanks! I wasnt sure if that was it or not...

jimthompson5910 (jim_thompson5910):

you're welcome